REGRESSION AND INVENTORIES

Jasper Furnishings has $350 million in sales. The company expects that its sales will increase 9% this year. Jasper's CFO uses a simple linear regression to forecast the company's inventory level for a given level of projected sales. On the basis of recent history, the estimated relationship between inventories and sales (in millions of dollars) is as follows:

Inventories = $35 + 0.155(Sales)

a. Given the estimated sales forecast and the estimated relationship between inventories and sales, what are your forecasts of the company's year-end inventory level? Enter your answer in millions. For example, an answer of $25,000,000 should be entered as 25. Round your answer to two decimal places.

b. What are your forecasts of the company's year-end inventory turnover ratio? Round your answer to two decimal places.
Answer #1

a). Sales Forecast = Sales + (1 + % increase in sales)

= $350 million * (1 + 0.09) = $381.50 million

Inventories = $35 million + 0.155(Sales)

= $35 million + 0.155($381.50 million)

= $35 million + $59.1325 million

= $94.1325 million, or $94.13 million

b). Inventory Turnover Ratio = Sales / Year-end Inventories

= 381.50 / 94.1325 = 4.05 times